OpenStudy (anonymous):

Simplify completely quantity 5 x squared plus 9 x minus 2 over quantity x squared plus 12 x plus 20 times quantity x squared plus 17 x plus 70 over quantity 15 x minus 3

OpenStudy (anonymous):

does it look like this? \[\frac{5x^2+9x-2}{x^2+12x+20} \times \frac{x^2+17+70}{15x-3}\]

OpenStudy (johnweldon1993):

\[\large \frac{5x^2+9x-2}{x^2+12x+20} \times \frac{x^2+17x+70}{15x-3}\] Just forgot an 'x' :)